Circular No. 5735

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOTICE TO SUBSCRIBERS As an experiment, the Central Bureau decided not to issue telegrams

(telexes, mailgrams) recently, specifically with regard to the comets and supernovae discovered during the past week. It has been felt for some time that the Circulars are getting such rapid and wide distribution electronically that the telegrams have become superfluous. There are now very few countries with active observatories that cannot receive electronic mail, and any amateur astronomer who subscribes to a widespread network, such as Compuserve, can also easily subscribe to the e-mail Circulars. It certainly seems not to be the case that the absence of telegrams in the past week has caused important observations to be missed. On the other hand, in comparison with the Circulars, the telegrams contain extremely little information, and they are also expensive both for subscribers to receive and for the Central Bureau to send (and to maintain address lists with the telegraph company). Comments from telegram subscribers are solicited. Unless there are good reasons for continuing the telegram service, the Central Bureau would like to phase it out, transferring any remaining credit balances to the corresponding Circulars accounts.

PERIODIC COMET FORBES (1993f)

Ephemeris from the orbital elements on MPC 16379:

1993 TT α_{2000} δ_{2000} Δ r ϵ β m_{1}
Mar. 24 21^{h}34.^{m}74 −18^{o}07.^{′}0 2.002 1.450 43.^{o}5 28.^{o}2 13.9
Apr. 3 22 05.21 −15 03.4 1.979 1.462 45.4 29.1 14.0
13 22 33.85 −11 51.4 1.960 1.481 47.5 30.0 14.0
23 23 00.72 − 8 36.4 1.943 1.507 49.9 30.7 14.1
May 3 23 25.85 − 5 23.0 1.926 1.540 52.6 31.4 14.2
13 23 49.34 − 2 14.9 1.909 1.579 55.7 31.9 14.4
23 0 11.22 + 0 45.0 1.890 1.622 59.1 32.4 14.5
June 2 0 31.50 + 3 34.8 1.867 1.671 62.9 32.7 14.7
12 0 50.17 + 6 13.1 1.840 1.723 67.2 32.9 14.9
22 1 07.15 + 8 39.0 1.808 1.778 71.9 32.9 15.0
July 2 1 22.28 +10 51.9 1.771 1.836 77.2 32.7 15.2
12 1 35.41 +12 51.7 1.730 1.896 83.0 32.1 15.4
22 1 46.27 +14 37.9 1.684 1.958 89.4 31.3 15.5
Aug. 1 1 54.57 +16 10.3 1.636 2.021 96.5 29.9 15.7

1993 March 30 Brian G. Marsden